PREDICTION (REGRESSION) – Chapter 12

3. A researcher was interested in whether there was a relationship between stress and depression scores obtained from emergency health care providers. The data from 10 emergency workers are below. This is the same data you calculated a correlation on in Question 1. We are trying to use stress scores (IV or Predictor) to predict depression scores (DV or outcome). Using SPSS, calculate a linear regression model:

Y = a+b1X1      where Y is Depression and X1 is stress.

We actually could rewrite the equation as:    Depression = a+b1Stress

Stress

Depression

38

26

28

16

42

34

18

22

26

15

45

24

32

18

26

18

22

12

33

15

Using the output:

  1. What is the value of b1? _______
    standard Error of b1? ___________

Significance Level of b1 (i.e. “p-value”)? ___________

  1. Write a sentence that describes the interpretation of b1.
  1. What is the standardized value of b1? Notice this is the same as the correlation coefficient from Question 1! The correlation coefficient and the standardized regression coefficient are the same in a one predictor regression.

  1. What is the value of a? _______
    standard Error of a? ___________

Significance Level of a (i.e. “p-value”)? ___________

  1. Write a sentence that describes the interpretation of a.
  1. What is the predicted value of Depression, for someone with a score of 30 on stress?
Answer #1

Answer:

Given that,

\rightarrow  A researcher was interested in whether there was a relationship between stress and depression scores obtained from emergency health care providers.

\rightarrow The data from 10 emergency workers are below. This is the same data you calculated a correlation on in Question 1.

\rightarrow We are trying to use stress scores (IV or Predictor) to predict depression scores (DV or outcome).

Calculate a linear regression model:

Y = a+b1X1      where Y is Depression and X1 is stress.

We actually could rewrite the equation as:    

Depression = a+b1Stress

Stress Depression
38 26
28 16
42 34
18 22
26 15
45 24
32 18
26 18
22 12
33 15

Using SPSS, go to Data, select Data analysis, Choose regression. Put stress in X input range and depression in Y input range.

Summary Output
Regression Statistics
Multiple R 0.647
R-square 0.418
Adjusted R square 0.345
Standard error 5.326
Observations 10
ANOVA
df SS MS F Significance F
Regression 1 163.072 163.072 5.749 0.043
Residual 8 226.928 28.366
Total 9 390.000
Coefficients Standard Error t stat P-value Lower 95% Upper 95%
Intercept 4.819 6.552 0.736 0.483 -10.289 19.927
Stress 0.490 0.204 2.398 0.043 0019 0.961

(a).

b1=0.49

Standard error of b1=0.204

The p-value of b1=0.043

(b).

With one unit increase in stress, depression increases by 0.49 units.

(c).

Standardized value of b1 (Multiple R)=0.647

(d).

a=4.819

Standard error of a=6.552

The p-value of a=0.483

(e).

If there is no stress, the depression score will be 4.819.

(f).

y=4.819+0.49x

Stress=30

Depression=4.819+0.49(30)

=4.819+14.7

=19.519